a three digit number is equal to17 times the sum of its digits.
if 198 is added to the number, the digits are interchanged.
the addition of first and third digit is 1less than middle digit.
find the number

Answer by Edwin McCravy(20055)

Let h = the hundreds digit
Let t = the tens digit
Let u = the units (or ones) digit

a three digit number is equal to 17 times the sum of its digits.
100h+10t+u = 17(h+t+u)
100h+10t+u = 17h+17t+17u
83h-7t-16u = 0

if 198 is added to the number, the digits are interchanged.
100h+10t+u+198 = 100u+10t+h
       99h-99u = -198
           h-u = -2

the addition of first and third digit is 1 less than middle digit.
  h+u = t-1
h-t+u = -1

find the number
Solve this system:

83h-7t-16u =  0
  h   -  u = -2
  h- t+  u = -1

You finish.

Edwin